find 3 rational numbers between 99999/99999  and  11999/99000

Grade:10

3 Answers

Prateek Pravanjan
8 years ago
I THINK WE CAN’T FIND......BECAUSE THESE TWO NUMBERS ARE THE SAME NUMBER
the first number=99999/99999=(9X11)999/99000X9999=9999X11999/99000X9999=11999/99000=THE SECOND NUMBER.........HENCE THERE CANT BE NO NUMBER BETWEEN THE TWO........HOPE U R SATISFIED WITH THE ANSWER
Abishek arun
8 years ago
No rational numbers between the two because the numbers are same 
     99999                    (11 × 9)999            11999 × 9999                                                 11999
 
= -----------   = --------------------  =  ----------------------  = 11999 × 9 -9000 =  --------
     99999                      99999                             99999                                                             99000
Same number, so rational number cant be found
Please view this in askiitians forum because it is clear here only, it is not clear in mail
please approve my anwer
Abishek arun
8 years ago
Sorry answer is not clear 
answer wont come because given numbers are same numbers